# Self-organization and spectral mechanism of attractor landscapes in high-capacity kernel Hopfield networks

### v6

@@ Abstract
Kernel-based learning methods can dramatically increase the storage capacity of Hopfield networks, yet the dynamical mechanism behind this enhancement remains poorly understood. We address this gap by unifying the geometric analysis of the attractor landscape with the spectral theory of kernel machines. Using a novel metric, Pinnacle Sharpness, we first uncover a rich phase diagram of attractor stability, identifying a Ridge of Optimization where the network achieves maximal robustness under high-load conditions. Phenomenologically, this ridge is characterized by a Force Antagonism, where a strong driving force is balanced by a collective feedback force. Theoretically, we reveal that this phenomenon arises from a specific reorganization of the weight spectrum, which we term Spectral Concentration. Unlike a simple rank-1 collapse, our analysis shows that the network on the ridge self-organizes into a critical state: the leading eigenvalue is amplified to maximize global stability (Direct Force), while the trailing eigenvalues are preserved to maintain high memory capacity (Indirect Force). These findings provide a complete physical picture of how high-capacity associative memories are formed, demonstrating that optimal performance is achieved by tuning the system to a spectral Goldilocks zone between rank collapse and diffusion.
@@ 1 Introduction
    Unified Physical Picture: We demonstrate that the observed Force Antagonism is the dynamical manifestation of Spectral Concentration. The dominant spectral mode creates a deep potential funnel (Direct Force), while the subordinate modes manage inter-pattern interference (Indirect Force).
@@ 3.2 Phenomenological Tool: Pinnacle Sharpness and Force Decomposition
To calculate the gradient, we treat the discrete state $\bm{s}$ as a continuous variable $\bm{x}\in\mathbb{R}^{N}$. Differentiating Eq. (3) yields the gradient vector, which we analytically decompose into two functionally distinct components:
@@ 5.1.1 Step 1: The Geometry-Spectrum Link
$M(\bm{\xi}^{\mu})\propto\lambda_{\max}(\bm{K}_{\alpha}).$ (9)
@@ 5.1.2 Step 2: Spectral Concentration on the Ridge
Proposition 2 (Spectral Concentration).
@@ 5.3 Unified Interpretation
The combination of phenomenological observation (Force Antagonism) and spectral evidence (Spectral Concentration) leads to a unified physical picture. The Ridge represents a sweet spot where the kernel locality $\gamma$ is tuned such that the system can amplify the principal spectral mode to form deep, sharp attractors (via Proposition 1), while effectively managing inter-pattern interference using the remaining spectral degrees of freedom. The Force Antagonism is the dynamical manifestation of this spectral disparity: the massive Direct Force arises from the dominant $\lambda_{1}$ component, while the Indirect Force reflects the collective contribution of the trailing modes, acting to refine the basin boundaries.
@@ 7 Conclusion
In this paper, we have presented a comprehensive theory of attractor self-organization in high-capacity kernel Hopfield networks. Moving beyond simple performance metrics, we uncovered the dynamical and spectral mechanisms that enable these networks to defy classical storage limits.
Our phenomenological analysis identified a Ridge of Optimization in the phase diagram, where the network achieves maximal stability characterized by a unique Force Antagonism. Crucially, our theoretical analysis revealed the spectral origin of this phenomenon: Spectral Concentration. We demonstrated that on the Ridge, the learning algorithm drives the weight spectrum into a critical state characterized by a massive leading eigenvalue for global stability, while retaining a non-zero heavy tail for high-capacity storage. This finding dispels the notion of a simple rank-1 collapse and instead portrays the optimal memory landscape as a finely tuned balance between attractive force and pattern discriminability.
By bridging the gap between the geometry of attractors and the spectrum of the kernel machine, this work establishes a solid foundation for designing next-generation associative memories. The principles of spectral concentration elucidated here offer a general blueprint for building robust, interpretable, and high-capacity memory systems, with potential implications extending to the analysis of modern attention-based architectures.
@@
We used Gemini 2.5 Pro for proofreading the English manuscript.

### v7

@@ Abstract
Kernel-based learning methods can dramatically increase the storage capacity of Hopfield networks, yet the dynamical mechanisms behind this enhancement remain poorly understood. We address this gap by combining a geometric characterization of the attractor landscape with the spectral theory of kernel machines. Using a novel metric, Pinnacle Sharpness, we empirically uncover a rich phase diagram of attractor stability, identifying a Ridge of Optimization where the network achieves maximal robustness under high-load conditions. Phenomenologically, this ridge is characterized by a Force Antagonism, in which a strong driving force is counterbalanced by a collective feedback force. We theoretically interpret this behavior as a consequence of a specific reorganization of the weight spectrum, which we term Spectral Concentration. Unlike a simple rank-1 collapse, our analysis shows that the network on the ridge self-organizes into a critical regime: the leading eigenvalue is amplified to enhance global stability (Direct Force), while the trailing eigenvalues remain finite to sustain high memory capacity (Indirect Force). Together, these results suggest a spectral mechanism by which learning reconciles stability and capacity in high-dimensional associative memory models.
@@ 1 Introduction
    Unified Physical Picture: We show that the observed Force Antagonism can be interpreted as a dynamical manifestation of Spectral Concentration. The dominant spectral mode creates a deep potential funnel (Direct Force), while the subordinate modes manage inter-pattern interference (Indirect Force).
@@ 3.2 Phenomenological Tool: Pinnacle Sharpness and Force Decomposition
To calculate the gradient, we treat the discrete state $\bm{s}$ as a continuous variable $\bm{x}\in\mathbb{R}^{N}$. This continuous relaxation is introduced solely as a local surrogate to probe the stability geometry in the neighborhood of discrete fixed points, rather than as an exact description of the underlying dynamics. Differentiating Eq. (3) yields the gradient vector, which we analytically decompose into two functionally distinct components:
@@ 3.3 Theoretical Framework: Spectral Representation
Information Geometric Framework: To analyze the learning dynamics, we also employ tools from Information Geometry. The KLR model forms a statistical manifold equipped with the Fisher Information Matrix (FIM) as a Riemannian metric. For a single neuron, the FIM $\bm{G}$ is given by $\bm{G}=\bm{K}\bm{D}\bm{K}$, where $\bm{D}$ is a diagonal matrix of prediction variances. This metric allows us to distinguish between the Euclidean Gradient $\nabla L$, which guides standard SGD, and the Natural Gradient $\tilde{\nabla}L=\bm{G}^{-1}\nabla L$ , which represents the steepest descent direction on the manifold. The interplay between these two gradients reveals the intrinsic geometry of the learning process. A detailed derivation of these concepts is provided in Appendix B.
@@ 5 Theoretical Analysis: Spectral Mechanism of Self-Organization
This section provides a theoretical interpretation of the phenomenology reported in Section 4.
@@ 5.1.1 Step 1: The Geometry-Spectrum Link
$M(\bm{\xi}^{\mu})\propto\lambda_{\max}(\bm{K}_{\alpha}),$ (9)
where the proportionality holds up to kernel-dependent constants under the local isometry approximation.
@@ 5.1.2 Step 2: Spectral Concentration on the Ridge
Proposition 2 (Spectral Concentration, empirical and heuristic).
@@ 5.3 Unified Interpretation
The combination of phenomenological observation (Force Antagonism) and spectral evidence (Spectral Concentration) leads to a unified physical picture. The Ridge represents a sweet spot where the kernel locality $\gamma$ is tuned such that the system can amplify the principal spectral mode to form deep, sharp attractors (via Proposition 1), while effectively managing inter-pattern interference using the remaining spectral degrees of freedom. The Force Antagonism can be interpreted as the dynamical manifestation of this spectral disparity: the massive Direct Force arises from the dominant $\lambda_{1}$ component, while the Indirect Force reflects the collective contribution of the trailing modes, acting to refine the basin boundaries.
@@ 6.1 The Ridge as a Critical Phase Boundary
6.2 The Information Geometry of Learning
Our spectral theory explains the static properties of the Ridge, but how the learning dynamics navigate to this critical state remains an open question. Here, we provide an information-geometric interpretation of the empirically observed learning dynamics, which suggests that the Ridge acts as a geometric barrier at the brink of singularity. As shown in Fig. B-1, the learning dynamics at the edge of the Ridge ($\gamma=0.01$) exhibit a pronounced divergence in the norm of the Natural Gradient. This behavior is consistent with an effective pinning of the learning trajectory near the boundary of the manifold.
Importantly, the approach to this regime is not chaotic but highly structured. As illustrated in Fig. B-2, the dynamics on the stable part of the Ridge ($\gamma=0.02$) display a two-phase learning process: an initial phase dominated by the acquisition of global stability, followed by a prolonged fine-tuning phase that refines memory capacity. From an information-geometric viewpoint, this behavior can be interpreted as learning guided by an emergent curved geometry. Consistent with this interpretation, the learning trajectory traces a distinct parabolic arc in parameter space (Fig. B-3), which is characteristic of gradient flow on a curved statistical manifold.
@@ 7 Conclusion
In this paper, we have presented a coherent theoretical account of attractor self-organization in high-capacity kernel Hopfield networks. Moving beyond conventional performance metrics, we investigated the dynamical and spectral mechanisms that allow these networks to operate near classical storage limits.
Our phenomenological analysis identified a Ridge of Optimization in the phase diagram, where the network achieves maximal stability and exhibits a distinctive Force Antagonism. We theoretically interpreted this behavior through its spectral origin, which we termed Spectral Concentration. On the Ridge, learning drives the weight spectrum toward a critical regime in which the leading eigenvalue is amplified to enhance global stability, while the remaining eigenvalues remain finite, enabling high-capacity storage. This picture goes beyond a simple rank-1 collapse and instead characterizes the optimal memory landscape as a finely balanced interplay between attractive forces and pattern discriminability.
By linking the geometry of attractors with the spectral structure of the kernel machine, this work provides a principled framework for understanding how learning induces self-organized memory states. More broadly, the notion of spectral concentration offers insight into how stability and capacity can be reconciled in associative systems, and may inform the analysis of related learning dynamics in modern kernel- and attention-based architectures.
@@
We used ChatGPT (GPT-5.2) and Gemini 2.5 Pro for proofreading the English manuscript.
@@ A.3 Theoretical Basis for Spectral Concentration (Proposition 2)
Appendix B Information Geometric Formulation
This appendix provides a brief overview of the information geometric concepts used in Section 6 to analyze the learning dynamics. We follow the standard formulation by Amari [19].
B.1 Statistical Manifold and Fisher Metric
The set of probability distributions defined by the KLR model forms a statistical manifold $M$. Each point on this manifold corresponds to a specific set of parameters $\bm{\alpha}$, which in information geometry are known as the natural parameters (or $\theta$-coordinates). The distance between two nearby points on this manifold, corresponding to parameters $\bm{\alpha}$ and $\bm{\alpha}+d\bm{\alpha}$, is measured by the infinitesimal change in the probability distribution. This distance is quantified by the Fisher Information Matrix (FIM) $\bm{G}(\bm{\alpha})$ , which acts as a Riemannian metric tensor.
For a single neuron, the probability of the output being +1+1 is given by the sigmoid of the logit $u=\sum_{u}\alpha_{u}K(\cdot,\bm{\xi}^{\mu})$. This follows a Bernoulli distribution, which is a member of the exponential family. For such models, the FIM is defined as the expectation of the outer product of the gradient of the log-likelihood:
$G_{\mu\nu}(\bm{\alpha})=\mathbb{E}\left[\frac{\partial\log p}{\partial\alpha_{\mu}}\frac{\partial\log p}{\partial\alpha_{\nu}}\right].$ (16)
A key property of the logistic function is that $\partial p/\partial\alpha_{\mu}=(y-p)K(\cdot,\bm{\xi}^{\mu})$, where $y$ is the target and $p$ is the model’s output probability. Taking the expectation over the data distribution (and assuming independence), this leads to the specific form used in the main text:
$\bm{G}=\bm{K}\bm{D}\bm{K},$ (17)
where $\bm{D}$ is a diagonal matrix with entries $D_{\mu\mu}=p_{\mu}(1-p_{\mu})$ representing the prediction variance for pattern $\bm{\xi}^{\mu}$.
B.2 Natural Gradient Descent
Standard gradient descent (SGD) follows the steepest descent direction in the Euclidean parameter space. However, this path is not necessarily the steepest on the curved statistical manifold. The direction of steepest descent on the manifold is given by the Natural Gradient [20].
The Natural Gradient $\tilde{\nabla}L$ is defined as the direction that maximizes the decrease in the loss function $L$ for a small, fixed “information distance” (measured by the KL-divergence). For an infinitesimal step, the KL-divergence between $p(\bm{\alpha})$ and $p(\bm{\alpha}+d\bm{\alpha})$ is approximated by the quadratic form of the Fisher metric:
$D_{\text{KL}}(p(\bm{\alpha})||p(\bm{\alpha}+d\bm{\alpha}))\approx\frac{1}{2}d\bm{\alpha}^{\top}\bm{G}(\bm{\alpha})d\bm{\alpha}.$ (18)
Minimizing the change in loss $\nabla L^{\top}d\bm{\alpha}$ subject to a fixed information distance leads to the Natural Gradient direction:
$\tilde{\nabla}L(\bm{\alpha})=\bm{G}(\bm{\alpha})^{-1}\nabla L(\bm{\alpha}).$ (19)
The Natural Gradient Descent (NGD) update rule is therefore $\bm{\alpha}_{t+1}=\bm{\alpha}_{t}-\eta\tilde{\nabla}L(\bm{\alpha}_{t})$. This formulation allows us to analyze the learning dynamics not as a simple descent in Euclidean space, but as a flow on a curved Riemannian manifold, revealing the geometric distortions induced by the FIM.
B.3 Experimental Analysis of Learning Dynamics
To validate the geometric interpretations presented in the Discussion, we conducted numerical analyses of the SGD learning trajectory. We tracked the parameter vector $\bm{\alpha}_{t}$ for a single representative neuron over 200 training steps under different hyperparameter regimes.
Gradient Norm Dynamics: We computed the squared norms of three distinct gradient vectors at each step $t$: the Euclidean norm $\|\nabla L(\bm{\alpha}_{t})\|^{2}$, the Riemannian norm $\|\nabla L(\bm{\alpha}_{t})\|^{2}_{G^{-1}}=\nabla L^{\top}G^{-1}\nabla L$, and and the Natural Step norm $\|G^{-1}\nabla L(\bm{\alpha}_{t})\|^{2}$. As shown in Fig. 6, the dynamics on the edge of the Ridge ($\gamma=0.01$) exhibit a divergence in the Riemannian and Natural Step norms. This indicates that the manifold is approaching a singularity, creating a geometric barrier for SGD.
Figure 6: Gradient Norm Dynamics on the Edge of the Ridge ($\gamma=0.01$). The divergence of the Riemannian and Natural Step norms indicates that the manifold is approaching a singularity, creating a geometric barrier for SGD.
Pythagorean Decomposition of Information Gain: Second, we decomposed the information gain at each step, $\Delta I_{t}\approx\frac{1}{2}\bm{\delta}_{t}^{T}G^{*}\bm{\delta}_{t}$ into two orthogonal components based on the eigenspace of the final FIM $G^{*}$: the Principal Gain (contribution from $\lambda_{1}$ and the Tail Gain (contribution from $\lambda_{k>1}$). As shown in Fig. 7, the dynamics on the stable part of the Ridge ($\gamma=0.02$) exhibit a two-phase process: an initial phase dominated by Principal Gain is followed by a prolonged Fine-tuning phase dominated by Tail Gain, demonstrating a hierarchical learning strategy.
Figure 7: Information Gain Decomposition on the Ridge ($\gamma=0.02$). The dynamics exhibit a two-phase process: an initial phase dominated by Principal Gain is followed by a prolonged Fine-tuning phase dominated by Tail Gain.
Geodesic Deviation: Finally, we analyzed the shape of the trajectory $\bm{\alpha}_{t}$ by plotting its deviation from the e-geodesic (straight line) against its progress towards the final state $\bm{\alpha}^{*}$. As shown in Fig. 8, the results reveal a stark contrast between the two regimes: while the Local trajectory ($\gamma=0.1$) is perfectly straight, the Ridge trajectory ($\gamma=0.02$) forms a distinct parabolic arc. This confirms that the trajectory on the Ridge is guided by the non-trivial curvature of the statistical manifold.
Figure 8: Geodesic Deviation of Learning Trajectories. The Ridge trajectory ($\gamma=0.02$, red solid line) exhibits a distinct parabolic arc, demonstrating that it is guided by the non-trivial curvature of the statistical manifold. In stark contrast, the Local trajectory ($\gamma=0.1$, blue dashed line) follows the e-geodesic (straight line) perfectly, indicating a flat, featureless learning landscape.